THE SPEED OF 5 KG BODY IS TO BE INCREASED FROM 10


M/S TO 20 M/S. WHAT CONSTANT FORCE IS NEEDED TO


PRODUCE THIS CHANGE IN SPEED IN EXACTLY 5 S ?

1
Expert's answer
2022-02-13T12:14:01-0500

Given:

"m=5\:\rm kg"

"v_1=10\;\rm m/s"

"v_2=20\;\rm m/s"

"\Delta t=5\:\rm s"


The Newton's second law says

"F=ma=m\frac{\Delta v}{\Delta t}"

"F=5*\frac{20-10}{5}=10\:\rm N"